# Prompts Are Programs Too! Understanding How Developers Build Software Containing Prompts

### v1

@@ Acknowledgements.
- Liu et al. (2024) Michael Xieyang Liu, Frederick Liu, Alexander J Fiannaca, Terry Koo, Lucas Dixon, Michael Terry, and Carrie J Cai. 2024. ”We need structured output”: Towards user-centered constraints on large language model output. In ACM CHI Conference on Human Factors in Computing Systems Extended Abstracts (CHI). 1–9. https://doi.org/10.1145/3613905.3650756
@@
- Liu et al. (2023) Pengfei Liu, Weizhe Yuan, Jinlan Fu, Zhengbao Jiang, Hiroaki Hayashi, and Graham Neubig. 2023. Pre-train, prompt, and predict: A systematic survey of prompting methods in natural language processing. ACM Computing Surveys (CSUR) 55, 9 (2023), 1–35. https://doi.org/10.1145/3560815
@@
- Ma et al. (2024) Wanqin Ma, Chenyang Yang, and Christian Kästner. 2024. (Why) is my prompt getting worse? Rethinking regression testing for evolving LLM APIs. In Proceedings of the IEEE/ACM 3rd International Conference on AI Engineering-Software Engineering for AI. 166–171.
@@
- McGhee et al. (2007) Gerry McGhee, Glenn R Marland, and Jacqueline Atkinson. 2007. Grounded theory research: Literature reviewing and reflexivity. Journal of Advanced Nursing 60, 3 (2007), 334–342. https://doi.org/10.1111/j.1365-2648.2007.04436.x
@@
- Mishra et al. (2023) Aditi Mishra, Utkarsh Soni, Anjana Arunkumar, Jinbin Huang, Bum Chul Kwon, and Chris Bryan. 2023. PromptAid: Prompt exploration, perturbation, testing and iteration using visual analytics for large language models. arXiv preprint arXiv:2304.01964 (2023).
@@
- Nahar et al. (2023) Nadia Nahar, Haoran Zhang, Grace Lewis, Shurui Zhou, and Christian Kästner. 2023. A meta-summary of challenges in building products with ML components–Collecting experiences from 4758+ practitioners. In IEEE/ACM International Conference on AI Engineering–Software Engineering for AI (CAIN). 171–183. https://doi.org/10.1109/CAIN58948.2023.00034
@@
- Nahar et al. (2022) Nadia Nahar, Shurui Zhou, Grace Lewis, and Christian Kästner. 2022. Collaboration challenges in building ml-enabled systems: Communication, documentation, engineering, and process. In IEEE/ACM International Conference on Software Engineering (ICSE)). 413–425. https://doi.org/10.1145/3510003.3510209
@@
- Patel et al. (2024) Harsh Patel, Dominique Boucher, Emad Fallahzadeh, Ahmed E Hassan, and Bram Adams. 2024. A State-of-the-practice release-readiness checklist for generative AI-based software products. arXiv preprint arXiv:2403.18958 (2024).
- Pei and Jurgens (2023) Jiaxin Pei and David Jurgens. 2023. When do annotator demographics matter? Measuring the influence of annotator demographics with the POPQUORN dataset. In Linguistic Annotation Workshop (LAW-XVII). Association for Computational Linguistics, 252–265. https://doi.org/10.18653/v1/2023.law-1.25
@@
- Qian et al. (2024) Crystal Qian, Emily Reif, and Minsuk Kahng. 2024. Understanding the dataset practitioners behind large language models. In ACM CHI Conference on Human Factors in Computing Systems Extended Abstracts (CHI). 1–7. https://doi.org/10.1145/3613905.3651007
- Santy et al. (2023) Sebastin Santy, Jenny Liang, Ronan Le Bras, Katharina Reinecke, and Maarten Sap. 2023. NLPositionality: Characterizing design biases of datasets and models. In Annual Meeting of the Association for Computational Linguistics (ACL). 9080–9102. https://doi.org/10.18653/v1/2023.acl-long.505
@@
- Sheil (1986) Beau Sheil. 1986. Datamation®: Power tools for programmers. In Readings in Artificial Intelligence and Software Engineering. Elsevier, 573–580. https://doi.org/10.1016/B978-0-934613-12-5.50048-3
- Staggers and Norcio (1993) Nancy Staggers and Anthony F. Norcio. 1993. Mental models: concepts for human-computer interaction research. International Journal of Man-machine studies 38, 4 (1993), 587–605.
- Strobelt et al. (2022) Hendrik Strobelt, Albert Webson, Victor Sanh, Benjamin Hoover, Johanna Beyer, Hanspeter Pfister, and Alexander M Rush. 2022. Interactive and visual prompt engineering for ad-hoc task adaptation with large language models. IEEE Transactions on Visualization and Computer Graphics 29, 1 (2022), 1146–1156. https://doi.org/10.1109/TVCG.2022.3209479
- Suri (2011) Harsh Suri. 2011. Purposeful sampling in qualitative research synthesis. Qualitative research journal 11, 2 (2011), 63–75.
@@
- Wan et al. (2019) Zhiyuan Wan, Xin Xia, David Lo, and Gail C Murphy. 2019. How does machine learning change software development practices? IEEE Transactions on Software Engineering 47, 9 (2019), 1857–1871. https://doi.org/10.1109/TSE.2019.2937083
@@
- White et al. (2024) Jules White, Sam Hays, Quchen Fu, Jesse Spencer-Smith, and Douglas C Schmidt. 2024. Chatgpt prompt patterns for improving code quality, refactoring, requirements elicitation, and software design. In Generative AI for Effective Software Development. Springer, 71–108. https://doi.org/10.1007/978-3-031-55642-5_4
- Wolf and Paine (2020) Christine T Wolf and Drew Paine. 2020. Sensemaking practices in the everyday work of AI/ML software engineering. In IEEE/ACM International Conference on Software Engineering Workshops. 86–92. https://doi.org/10.1109/TSE.2019.2937083
- Wu et al. (2022a) Tongshuang Wu, Ellen Jiang, Aaron Donsbach, Jeff Gray, Alejandra Molina, Michael Terry, and Carrie J Cai. 2022a. Promptchainer: Chaining large language model prompts through visual programming. In ACM CHI Conference on Human Factors in Computing Systems Extended Abstracts (CHI). 1–10. https://doi.org/10.1145/3491101.3519729
- Wu et al. (2022b) Tongshuang Wu, Michael Terry, and Carrie Jun Cai. 2022b. Ai chains: Transparent and controllable human-ai interaction by chaining large language model prompts. In ACM CHI Conference on Human Factors in Computing Systems (CHI). 1–22. https://doi.org/10.1145/3491102.3517582
- Ye et al. (2024) Qinyuan Ye, Mohamed Ahmed, Reid Pryzant, and Fereshte Khani. 2024. Prompt engineering a prompt engineer. In Findings of the Association for Computational Linguistics ACL (ACL), Lun-Wei Ku, Andre Martins, and Vivek Srikumar (Eds.). 355–385.
@@
- Zamfirescu-Pereira et al. (2023b) JD Zamfirescu-Pereira, Heather Wei, Amy Xiao, Kitty Gu, Grace Jung, Matthew G Lee, Bjoern Hartmann, and Qian Yang. 2023b. Herding AI cats: Lessons from designing a chatbot by prompting GPT-3. In ACM Designing Interactive Systems Conference (DIS). 2206–2220. https://doi.org/10.1145/3563657.3596138
- Zamfirescu-Pereira et al. (2023c) JD Zamfirescu-Pereira, Richmond Y Wong, Bjoern Hartmann, and Qian Yang. 2023c. Why Johnny can’t prompt: How non-AI experts try (and fail) to design LLM prompts. In ACM CHI Conference on Human Factors in Computing Systems (CHI). 1–21. https://doi.org/10.1145/3544548.3581388
- Ziegler et al. (2022) Albert Ziegler, Eirini Kalliamvakou, X Alice Li, Andrew Rice, Devon Rifkin, Shawn Simister, Ganesh Sittampalam, and Edward Aftandilian. 2022. Productivity assessment of neural code completion. In ACM SIGPLAN International Symposium on Machine Programming (MAPS). 21–29.
- Ziegler et al. (2024) Albert Ziegler, Eirini Kalliamvakou, X Alice Li, Andrew Rice, Devon Rifkin, Shawn Simister, Ganesh Sittampalam, and Edward Aftandilian. 2024. Measuring GitHub Copilot’s impact on productivity. Communications of the ACM (CACM) 67, 3 (2024), 54–63. https://doi.org/10.1145/3633453

### v2

@@ Acknowledgements.
- Liu et al. (2024) Michael Xieyang Liu, Frederick Liu, Alexander J Fiannaca, Terry Koo, Lucas Dixon, Michael Terry, and Carrie J Cai. 2024. ”We need structured output”: Towards user-centered constraints on large language model output. In ACM CHI Conference on Human Factors in Computing Systems Extended Abstracts (CHI). 1–9. doi:10.1145/3613905.3650756
@@
- Liu et al. (2023) Pengfei Liu, Weizhe Yuan, Jinlan Fu, Zhengbao Jiang, Hiroaki Hayashi, and Graham Neubig. 2023. Pre-train, prompt, and predict: A systematic survey of prompting methods in natural language processing. ACM Computing Surveys (CSUR) 55, 9 (2023), 1–35. doi:10.1145/3560815
- Luo et al. (2014) Qingzhou Luo, Farah Hariri, Lamyaa Eloussi, and Darko Marinov. 2014. An empirical analysis of flaky tests. In ACM SIGSOFT International Symposium on Foundations of Software Engineering (FSE). 643–653. doi:10.1145/2635868.2635920
@@
- Ma et al. (2024) Wanqin Ma, Chenyang Yang, and Christian Kästner. 2024. (Why) is my prompt getting worse? Rethinking regression testing for evolving LLM APIs. In Proceedings of the IEEE/ACM 3rd International Conference on AI Engineering-Software Engineering for AI. 166–171. doi:10.1145/3644815.3644950
- Maalej et al. (2014) Walid Maalej, Rebecca Tiarks, Tobias Roehm, and Rainer Koschke. 2014. On the comprehension of program comprehension. ACM Transactions on Software Engineering and Methodology (TOSEM) 23, 4 (2014), 1–37. doi:10.1145/26226
@@
- McGhee et al. (2007) Gerry McGhee, Glenn R Marland, and Jacqueline Atkinson. 2007. Grounded theory research: Literature reviewing and reflexivity. Journal of Advanced Nursing 60, 3 (2007), 334–342. doi:10.1111/j.1365-2648.2007.04436.x
@@
- Mishra et al. (2025) Aditi Mishra, Bretho Danzy, Utkarsh Soni, Anjana Arunkumar, Jinbin Huang, Bum Chul Kwon, and Chris Bryan. 2025. PromptAid: Visual prompt exploration, perturbation, testing and iteration for large language models. IEEE Transactions on Visualization and Computer Graphics (2025). doi:10.1109/TVCG.2025.3535332
@@
- Nahar et al. (2023) Nadia Nahar, Haoran Zhang, Grace Lewis, Shurui Zhou, and Christian Kästner. 2023. A meta-summary of challenges in building products with ML components–Collecting experiences from 4758+ practitioners. In IEEE/ACM International Conference on AI Engineering–Software Engineering for AI (CAIN). 171–183. doi:10.1109/CAIN58948.2023.00034
@@
- Nahar et al. (2022) Nadia Nahar, Shurui Zhou, Grace Lewis, and Christian Kästner. 2022. Collaboration challenges in building ml-enabled systems: Communication, documentation, engineering, and process. In IEEE/ACM International Conference on Software Engineering (ICSE). 413–425. doi:10.1145/3510003.3510209
@@
- Patel et al. (2025) Harsh Patel, Dominique Boucher, Emad Fallahzadeh, Ahmed E Hassan, and Bram Adams. 2025. A state-of-the-practice release-readiness checklist for generative AI-based software products: A gray literature survey. IEEE Software 42, 01 (2025), 74–83. doi:10.1109/MS.2024.3440190
- Pei and Jurgens (2023) Jiaxin Pei and David Jurgens. 2023. When do annotator demographics matter? Measuring the influence of annotator demographics with the POPQUORN dataset. In Linguistic Annotation Workshop (LAW-XVII). Association for Computational Linguistics, 252–265. doi:10.18653/v1/2023.law-1.25
@@
- Qian et al. (2024) Crystal Qian, Emily Reif, and Minsuk Kahng. 2024. Understanding the dataset practitioners behind large language models. In ACM CHI Conference on Human Factors in Computing Systems Extended Abstracts (CHI). 1–7. doi:10.1145/3613905.3651007
- Santy* et al. (2023) Sebastin Santy*, Jenny T Liang*, Ronan Le Bras, Katharina Reinecke, and Maarten Sap. 2023. NLPositionality: Characterizing design biases of datasets and models. In Annual Meeting of the Association for Computational Linguistics (ACL). 9080–9102. doi:10.18653/v1/2023.acl-long.505
@@
- Sheil (1986) Beau Sheil. 1986. Datamation®: Power tools for programmers. In Readings in Artificial Intelligence and Software Engineering. Elsevier, 573–580. doi:10.1016/B978-0-934613-12-5.50048-3
- Staggers and Norcio (1993) Nancy Staggers and Anthony F Norcio. 1993. Mental models: concepts for human-computer interaction research. International Journal of Man-machine studies 38, 4 (1993), 587–605. doi:10.1006/imms.1993.1028
- Strobelt et al. (2022) Hendrik Strobelt, Albert Webson, Victor Sanh, Benjamin Hoover, Johanna Beyer, Hanspeter Pfister, and Alexander M Rush. 2022. Interactive and visual prompt engineering for ad-hoc task adaptation with large language models. IEEE Transactions on Visualization and Computer Graphics 29, 1 (2022), 1146–1156. doi:10.1109/TVCG.2022.3209479
- Suri (2011) Harsh Suri. 2011. Purposeful sampling in qualitative research synthesis. Qualitative research journal 11, 2 (2011), 63–75. doi:10.3316/QRJ1102063
@@
- Wan et al. (2019) Zhiyuan Wan, Xin Xia, David Lo, and Gail C Murphy. 2019. How does machine learning change software development practices? IEEE Transactions on Software Engineering 47, 9 (2019), 1857–1871. doi:10.1109/TSE.2019.2937083
@@
- White et al. (2024) Jules White, Sam Hays, Quchen Fu, Jesse Spencer-Smith, and Douglas C Schmidt. 2024. Chatgpt prompt patterns for improving code quality, refactoring, requirements elicitation, and software design. In Generative AI for Effective Software Development. Springer, 71–108. doi:10.1007/978-3-031-55642-5_4
- Wolf and Paine (2020) Christine T Wolf and Drew Paine. 2020. Sensemaking practices in the everyday work of AI/ML software engineering. In IEEE/ACM International Conference on Software Engineering Workshops. 86–92. doi:10.1109/TSE.2019.2937083
- Wu et al. (2022a) Tongshuang Wu, Ellen Jiang, Aaron Donsbach, Jeff Gray, Alejandra Molina, Michael Terry, and Carrie J Cai. 2022a. Promptchainer: Chaining large language model prompts through visual programming. In ACM CHI Conference on Human Factors in Computing Systems Extended Abstracts (CHI). 1–10. doi:10.1145/3491101.3519729
- Wu et al. (2022b) Tongshuang Wu, Michael Terry, and Carrie Jun Cai. 2022b. Ai chains: Transparent and controllable human-ai interaction by chaining large language model prompts. In ACM CHI Conference on Human Factors in Computing Systems (CHI). 1–22. doi:10.1145/3491102.3517582
- Ye et al. (2024) Qinyuan Ye, Mohamed Ahmed, Reid Pryzant, and Fereshte Khani. 2024. Prompt engineering a prompt engineer. In Findings of the Association for Computational Linguistics ACL (ACL), Lun-Wei Ku, Andre Martins, and Vivek Srikumar (Eds.). 355–385. doi:10.18653/v1/2024.findings-acl.21
@@
- Zamfirescu-Pereira et al. (2023b) JD Zamfirescu-Pereira, Heather Wei, Amy Xiao, Kitty Gu, Grace Jung, Matthew G Lee, Bjoern Hartmann, and Qian Yang. 2023b. Herding AI cats: Lessons from designing a chatbot by prompting GPT-3. In ACM Designing Interactive Systems Conference (DIS). 2206–2220. doi:10.1145/3563657.3596138
- Zamfirescu-Pereira et al. (2023c) JD Zamfirescu-Pereira, Richmond Y Wong, Bjoern Hartmann, and Qian Yang. 2023c. Why Johnny can’t prompt: How non-AI experts try (and fail) to design LLM prompts. In ACM CHI Conference on Human Factors in Computing Systems (CHI). 1–21. doi:10.1145/3544548.3581388
- Ziegler et al. (2022) Albert Ziegler, Eirini Kalliamvakou, X Alice Li, Andrew Rice, Devon Rifkin, Shawn Simister, Ganesh Sittampalam, and Edward Aftandilian. 2022. Productivity assessment of neural code completion. In ACM SIGPLAN International Symposium on Machine Programming (MAPS). 21–29. doi:10.1145/3520312.3534864
- Ziegler et al. (2024) Albert Ziegler, Eirini Kalliamvakou, X Alice Li, Andrew Rice, Devon Rifkin, Shawn Simister, Ganesh Sittampalam, and Edward Aftandilian. 2024. Measuring GitHub Copilot’s impact on productivity. Communications of the ACM (CACM) 67, 3 (2024), 54–63. doi:10.1145/3633453